OpenStudy (anonymous):

-Apply Sets to number and Functions- Let U be the set of integers from 0 to 10. Find (a) A U B and A ∩ B , (b) A and B, and (c) A x B for the Specified sets A and B 1.) A= {1, 3,5, 7, 9} and B= {3,6,9} 2.) A= {1,2, 3, 4, ,5 ,6 } and B={4, ,5 ,6 ,7 ,8} 3.) A={0,2,4,6,8,10} And B= {1,3,5,7,9} 4.) A= {0,5,10} and B= {1,4,7,10} Could you explain how you got answers? im not sure how to solve these

OpenStudy (across):

\[A=\left\{1,3,5,7,9\right\},\]\[B=\left\{3,6,9\right\}.\] \[A\cup B=\left\{1,3,5,6,7,9\right\},\]\[A\cap B=\left\{3,9\right\}.\] What do you mean with "A and B"?

OpenStudy (anonymous):

I dunno i just copied what was in my Text book :/ ill check for you

OpenStudy (anonymous):

are those answers for Questions 1 and 2 ?

OpenStudy (anonymous):

it doesnt tell me what A and b means lol, says somthing about Union and InterSection? if im not mistaken

OpenStudy (anonymous):

does that mean anythin to you?

OpenStudy (across):

A and B are sets. The union of two sets A and B is always the "sum," so to speak, of all their elements where those who repeat are left out. For example, A = {1,3,5,7,9}, B = {3,6,9}, A∪B = {1,3,5,6,7,9} The intersection of two sets A and B is always the elements common to both sets A and B. That is, only the elements which are in both sets are in their intersection. For example, A = {1,3,5,7,9}, B = {3,6,9}, A∩B = {3,9}. The cartesian product of two sets A and B is the total number of combinations in between the elements of both sets. For example, A = {1,3,5,7,9}, B = {3,6,9}, A×B = {(1,3),(1,6),(1,9),(3,3),(3,6),(3,9),(5,3),(5,6),(5,9),(7,3),(7,6),(7,9),(9,3),(9,6),(9,9)}

OpenStudy (anonymous):

could you # them i dont kno what those are are to whcihc of my Questions? so sorry :/

OpenStudy (across):

What I just wrote up there is the general procedure to solve all of these problems.

OpenStudy (across):

When is this due? I need to head out in a little, but I will be back soon.

OpenStudy (anonymous):

tommarrow ^^, but ill be on for 5 ish ours on FB so go out and Parrty ill be on for a while

OpenStudy (anonymous):

haha ^